OpenStudy (anonymous):

Differential Equation. e^(x)ydy/dx=e^(-y)+e^(-2x-y) I can't seem to figure this out.

OpenStudy (anonymous):

Is it: \[e^xy\frac{dy}{dx}=e^{-y}+e^{-2x-y}\] ?

OpenStudy (anonymous):

Yes that is correct

OpenStudy (anonymous):

\[\eqalign{ &e^xy\frac{dy}{dx}=e^{-y}+e^{-2x-y} \\ &e^xy\frac{dy}{dx}=e^{-y}+e^{-2x}e^{-y} \\ &e^xy\frac{dy}{dx}=e^{-y}(1+e^{-2x}) \\ &\frac{y}{e^{-y}}\frac{dy}{dx}=\frac{1+e^{-2x}}{e^x} \\ &y(e)^ydy=\frac{1+e^{-2x}}{e^x}dx \\ &y(e)^ydy=e^{-x}+e^{-3x}dx \\ &\int y(e)^ydy=\int e^{-x}+e^{-3x}dx \\ &\phantom{.......}LS=RS }\] Let's compute each side

OpenStudy (anonymous):

\[LS=\int ye^ydy=\int u\phantom{.}dv\] \[\eqalign{ &u=y\to du=(1) \\ &dv=e^y\to v=e^y \\ }\] \[\eqalign{ LS=\int ye^ydy=\int u\phantom{.}dv&=uv-\int v\phantom{.}du \\ &=ye^y-\int e^y \\ &=ye^y-e^y }\]

OpenStudy (anonymous):

\[RS=\int e^{-x}+e^{-3x}dx=-e^{-x}-\frac{1}{3}e^{-3x}+c_1\]

OpenStudy (anonymous):

Therefore, we obtain: \[ye^y-e^y=-e^{-x}-\frac{1}{3}e^{-3x}+c_1\]

OpenStudy (anonymous):

Thank you very much. I just had a problem finding what method to use.

OpenStudy (anonymous):

You would try and separate the ODE :)